# ReorderBench: A Benchmark for Matrix Reordering

### v1

@@ V-C Formal Study
In this study, the experts answer the questions independently. They achieve an ICC of 0.802 with a 95% confidence interval of [0.786, 0.817]. This level of agreement ensures the reliability of expert assessments for further analysis.
To compare the quality metrics with expert assessments, we first rank the variations according to the comparison results of each matrix. One issue in assigning the ranks is that, despite the substantial agreement among the experts, several conflicts still arise regarding the comparison results. To address this issue, we employ the Elo rating system [41] to handle such conflicts. This system adjusts the ratings of the variations to align with the comparison results, thereby reflecting the quality judged by the experts. The variation with the highest rating is ranked $1$, with subsequent ranks assigned in descending order of rating. After assigning the ranks, we use Kendall’s coefficient of rank correlation to measure the level of consistency between each quality metric and human experts.
Fig. 13 shows the distribution of the rank correlation on binary and continuous matrices. Following the common practice in empirical studies [42, 43], we perform Friedman tests to compare the rank correlation of the quality metrics. The results show that the correlation difference among the five quality metrics is significant in both binary matrices ($\chi^{2}(4)=61.96$, $p<0.001$) and continuous matrices ($\chi^{2}(4)=62.23$, $p<0.001$). The pairwise Wilcoxon signed-rank test further indicates that the convolution- and entropy-based scoring method significantly outperforms existing metrics ($p<0.05$). These results demonstrate that our scoring method aligns with the experts in assessing the visual pattern quality.
@@ VI Benchmark Applications and Analysis
In this section, we illustrate the potential utility of our benchmark on three applications. First, it enables the evaluation of reordering algorithms by comparing their performance on the benchmark. Second, it supports the development of a deep scoring model for measuring the quality of visual patterns. Third, it enables the development of a matrix reordering model based on the metric provided by the scoring model.
@@ VI-A Evaluating Reordering Algorithms
We compare $44$ reordering algorithms on the ReorderBench test set to assess their effectiveness in revealing visual patterns.
TABLE II: Evaluated base matrix reordering algorithms.
@@
Graph OPTICS [52], RCM [53], SPIN [54], TSP [55], VAT [56]
Reordering algorithms. Behrisch et al.  [20] classify existing automated reordering algorithms into 6 categories: Robinsonian, spectral, dimension reduction, heuristic, graph-theoretic, and biclustering. We select commonly used reordering algorithms in each category except for biclustering, as they are not suitable for symmetric matrices. This results in 44 algorithms, including 18 base algorithms and their variations. The selected base algorithms are listed in Table II. The full list of evaluated algorithms is available in the supplemental material. The implementations of 39 algorithms are sourced from the R package seriation [32]. We use the RCM algorithm from the Python SciPy library [57] and implement 4 algorithms in the heuristic category in Python for which we cannot find any reference implementation.
@@
TABLE III: Evaluation results of existing matrix reordering algorithms and our matrix reordering model described in Sec. VI-C. The best one is bold, and the runner-up is underlined.
@@
Results and analysis. We evaluate existing reordering algorithms on matrices of all four sizes. To identify the most effective algorithms within each algorithm category, we apply a selection process based on the performance averaged over matrix sizes. The selection favors algorithms with the best performance on individual patterns and those that rank as the top two in performance further averaged over all patterns. In our case, this results in two selected algorithms per category since the top-performing algorithm on individual patterns overlaps with one of the top two on average performance across patterns. Table III shows the average performance of these algorithms. The full results are available in the supplemental material. Both the average results and individual results on each size demonstrate that existing algorithms perform well on block and off-diagonal block patterns but poorly on star and band patterns. This points to the research potential for developing algorithms that can effectively reveal star and band patterns. The two algorithms that achieve the best performance for block, off-diagonal block, and star patterns are OLO_ward and SPIN. OLO_ward, a variation of the optimal leaf ordering algorithm [1], refines the hierarchical clustering dendrogram obtained with Ward’s linkage [59]. Hierarchical clustering captures the global structure of the matrix. Meanwhile, the optimal leaf ordering algorithm optimizes the distance between adjacent leaves, which further performs local optimization to reveal visual patterns. SPIN, the sorting points into neighborhoods algorithm [54], iteratively relocates each row to its most suitable neighborhood. Inspired by simulated annealing, the method starts by exploring large-scale relocations to build the global layout, then gradually reduces the relocation scale to capture the local structure. The commonality of the two methods in considering both global and local structures highlights the importance of integrating these two perspectives to reveal visual patterns. Although Barycenter achieves the best performance for band patterns, its average performance score ($0.366$) is considerably lower than OLO_ward ($0.573$) and SPIN ($0.542$) and ranks sixth-to-last among the $44$ evaluated algorithms. Therefore, we exclude it from our discussion.
@@
VI-B Building a Deep Scoring Model
The accuracy of our scoring method in measuring the quality of visual patterns makes it an appealing optimization criterion for reordering algorithms. However, this method requires prior knowledge of the types and sizes of the visual patterns to configure the convolutional kernels. This prerequisite, often absent in matrices beyond this benchmark, limits the generalizability of the scoring method. To address this issue, we build a unified scoring model based on ReorderBench. This model aligns with the convolution- and entropy-based scoring method across all four visual patterns in both binary and continuous matrices and can also measure matrices of varying sizes. As shown in Fig. 14, the key feature of this model is that it does not require prior knowledge of the patterns in the matrix, which greatly improves its generalizability. Next, we briefly introduce how to build this model.
@@ VI-B Building a Deep Scoring Model
Model training. Each ReorderBench sample includes: 1. matrix $A_{i}$, 2. the type of visual pattern in its template $pt_{i}$, 3. the ground-truth quality score $S_{i}$, and 4. the number of index swaps applied to the matrix $sw_{i}$. A key characteristic of matrices with high-quality visual patterns is that similar rows are adjacent. Existing scoring methods check this adjacency by taking the dissimilarity matrix as input [32]. Following this practice, we augment each sample with dissimilarity matrix $D_{i}$, which is derived from $A_{i}$ using the Euclidean distance between its rows.
@@
Scoring matrices of varying sizes. To predict scores for matrices of varying sizes, the key is transforming the matrix to the desired size while preserving visual patterns. Previous research in computer vision has shown that resizing an image effectively achieves this goal and incorporates it as a routine step in image pre-processing [65]. Therefore, to score matrices of different sizes, we convert them to images and resize them to the training size of the model. The score is then predicted by the model based on the resized images. During training and evaluation, we resize all matrices to $200{\mkern-2.0mu\times\mkern-2.0mu}200$.
@@
We evaluate the three scoring models on matrices of all four sizes. Table IV shows their average performance. The full results are available in the supplemental material. On all matrix sizes, ConvNeXt-T consistently performs the best in terms of both accuracy and efficiency. Therefore, we employ ConvNeXt-T to build the unified scoring model. We demonstrate the capability of the scoring model to handle real-world matrices by applying it to the Pajek graph collection [8] and the Network Repository [12]. To align with the training samples, we normalize these matrices so that their entries fall within $[0,1]$. As shown in Fig. 15, the unified scoring model generates reasonable scores for these matrices.
@@
VI-C Building a Matrix Reordering Model
The representative and diverse matrices in ReorderBench offer valuable supervision for training a deep reordering model. By treating the matrices with index swaps as negative samples and their ground-truth matrices as positive samples, we build a deep model for matrix reordering based on ResNet-18 [60] due to its demonstrated performance. The model architecture is introduced in the supplemental material.
@@ VI-C Building a Matrix Reordering Model
We train eight deep reordering models for each matrix size, with four dedicated to binary matrices and four to continuous matrices. Within each matrix type, four models correspond to four different visual patterns. All these models are trained with a batch size of 512 using AdamW optimizer with a base learning rate of 0.01. For each model, we train for 120 epochs and employ a cosine annealing scheduler with 20 epochs of linear warm-up. To further enhance the model performance and robustness, we perform test-time augmentation based on the unified scoring model. More details on the augmentation are described in the supplemental material.
Results. The last row of Table III presents the performance of our deep reordering model averaged over all matrix sizes. The full results are available in the supplemental material. For the block pattern, our model performs slightly better than existing methods. As a central topic for matrix reordering research, the task of revealing block patterns is well studied and thus poses a challenge for further improvement. For the off-diagonal block, star, and band patterns, our deep reordering model achieves higher performance scores than existing reordering methods. Although ResNet-18 may seem basic by current standards, the observed improvement confirms the value of our benchmark. Moreover, although reordering continuous matrices is generally harder than binary matrices, our model achieves higher performance scores for all four visual patterns in continuous matrices due to the larger amount of training data. This also highlights the importance of large-scale datasets for training high-performance deep reordering models.
@@ VII Discussion and Future Work
Support hybrid patterns. Although the proposed pipeline achieves impressive results in generating matrices with the four visual patterns, the creation of matrices with hybrid patterns, such as the combination of block + band, poses two challenges. The first is how to generate representative matrices with hybrid patterns. Addressing this involves exploring methods to integrate different patterns into hybrid ones, which deserves further investigation. A promising method is to adapt the mix-up technique [66], commonly used to augment training data. This technique improves the robustness and generalization of the model by blending multiple images and their labels to create new composite samples. The second lies in assessing hybrid patterns. As different types of visual patterns in a matrix can potentially obscure each other, and new patterns may appear during variation generation, it is difficult to measure the quality of these hybrid patterns by simply combining their scores derived by our scoring method. For example, Fig.  16(a) shows a matrix with a hybrid pattern of off-diagonal block and band. After 30 times of index swaps, these two visual patterns obscure each other, and many star patterns also appear (Fig.  16(b)). This highlights the importance of exploring effective methods to identify and measure the quality of hybrid patterns when they are mixed during the variation process.
@@
Reorder matrices of varying sizes with one model. Our evaluation demonstrates that, compared with existing algorithms, deep models greatly improve reordering performance by predicting the best permutation instead of performing a step-by-step search. However, the adaptability of deep reordering models to matrices with varying sizes remains an issue. These models are generally designed for and trained on matrices of fixed size [4, 3]. This highlights a potential area for innovation, enabling these models to accommodate matrices of different sizes. For smaller matrices, padding the matrix with empty rows and columns allows it to be reordered by the deep model. However, trials on our reordering model show that it occasionally ignores the matrix with smaller patterns, as the addition of many empty rows and columns makes these patterns nearly indistinguishable from noise. Thus, developing techniques to preserve original patterns during padding warrants further exploration. For larger matrices, one possible solution is to employ the divide-and-conquer strategy. The matrix is first divided into smaller matrices, each of which is reordered independently. Then they are combined into the final reordered matrix. However, how to divide the matrix without damaging the visual patterns remains an issue, which deserves further exploration.
@@ VIII Conclusion
In this paper, we introduce ReorderBench, a benchmark designed to evaluate and improve matrix reordering techniques. It features a representative and diverse collection of binary and continuous matrices, and incorporates a convolution- and entropy-based scoring method. Three applications have demonstrated that this benchmark not only serves as a resource for evaluating different reordering algorithms but can also enable more effective and efficient development of a unified scoring model and deep reordering models. These improvements enhance our ability to reveal and interpret the underlying structures within complex datasets.

### v2

@@ V-C Formal Study
In this study, the experts answer the questions independently. They achieve an ICC of 0.789 with a 95% confidence interval of [0.772, 0.806]. This level of agreement ensures the reliability of expert assessments for further analysis.
To compare the quality metrics with expert assessments, we first rank the variations according to the comparison results of each matrix. One issue in assigning the ranks is that, despite the substantial agreement among the experts, several conflicts still arise regarding the comparison results. To address this, we employ the Elo rating system [51] to handle such conflicts. This system adjusts the ratings of the variations to align with the comparison results, thereby reflecting the quality judged by the experts. The variation with the highest rating is ranked $1$, with subsequent ranks assigned in descending order of rating. After assigning the ranks, we use Kendall’s coefficient of rank correlation to measure the level of consistency between each quality metric and human experts.
Fig. 13 shows the distribution of the rank correlation on binary and continuous matrices. Following the common practice in empirical studies [52, 53], we perform Friedman tests to compare the rank correlation of the quality metrics. The results show that the correlation difference among the five quality metrics is significant in both binary matrices ($\chi^{2}(4)=$ $56.02$, $p<0.001$) and continuous matrices ($\chi^{2}(4)=$ $64.51$, $p<0.001$). The pairwise Wilcoxon signed-rank test further indicates that the convolution- and entropy-based scoring method significantly outperforms existing metrics ($p<0.05$). These results demonstrate that our scoring method aligns with the experts in assessing the visual pattern quality.
@@ VI Benchmark Applications and Analysis
In this section, we illustrate the potential utility of our benchmark on three applications. First, it enables the evaluation of reordering algorithms by comparing their performance on the benchmark. Second, it supports the development of a deep scoring model for measuring the quality of visual patterns. Third, it enables the development of a matrix-reordering model based on the metric provided by the scoring model. The evaluations presented here are conducted on the generated matrices. For those on the real-world matrices, please refer to Sec. 9.4 of the supplemental material.
@@ VI-A Evaluating Reordering Algorithms
We compare $45$ reordering algorithms on the ReorderBench test set to assess their effectiveness in revealing visual patterns.
TABLE II: Evaluated base matrix-reordering algorithms.
@@
Graph NN__\__2OPT [21], OPTICS [62], RCM [63],
SPIN [64], TSP [65], VAT [66]
Reordering algorithms. Behrisch et al.  [20] classify existing automated reordering algorithms into 6 categories: Robinsonian, spectral, dimension reduction, heuristic, graph-theoretic, and biclustering. We select commonly used reordering algorithms in each category except for biclustering, as they are not suitable for symmetric matrices. This results in 45 algorithms, including 19 base algorithms and their variations. The selected base algorithms are listed in Table II. The full list of evaluated algorithms is available in Table 2 of the supplemental material. The implementations of 39 algorithms are sourced from the R package seriation [39]. We use the RCM algorithm from the Python SciPy library [67], the NN__\__2OPT algorithm from the Reorder.js library [68], and implement 4 algorithms in the heuristic category in Python for which we cannot find any reference implementation.
@@
TABLE III: Evaluation results of existing matrix-reordering algorithms and our matrix-reordering model described in Sec. VI-C. The best one is bold, and the runner-up is underlined.
@@
Results and analysis. We evaluate existing reordering algorithms on matrices of all four sizes. To identify the most effective algorithms within each algorithm category, we apply a selection process based on the performance averaged over matrix sizes. The selection favors algorithms with the best performance on individual patterns and those that rank as the top two in performance further averaged over all patterns. In our case, this results in two selected algorithms per category since the top-performing algorithm on individual patterns overlaps with one of the top two on average performance across patterns. Table III shows the average performance of these algorithms. The full results are available in Sec. 11 of the supplemental material. Both the average results and individual results on each size demonstrate that existing algorithms perform well on block and off-diagonal block patterns but poorly on star and band patterns. This points to the research potential for developing algorithms that can effectively reveal star and band patterns. The two algorithms that achieve the best performance for block, off-diagonal block, and star patterns are OLO_ward and SPIN. OLO_ward, a variation of the optimal leaf ordering algorithm [1], refines the hierarchical clustering dendrogram obtained with Ward’s linkage [70]. Hierarchical clustering captures the global structure of the matrix. Meanwhile, the optimal leaf ordering algorithm optimizes the distance between adjacent leaves, which further performs local optimization to reveal visual patterns. SPIN, the sorting points into neighborhoods algorithm [64], iteratively relocates each row to its most suitable neighborhood. Inspired by simulated annealing, the method starts by exploring large-scale relocations to build the global layout, then gradually reduces the relocation scale to capture the local structure. The commonality of the two methods in considering both global and local structures highlights the importance of integrating these two perspectives to reveal visual patterns. Although Barycenter achieves the best performance for band patterns, its average performance score ($0.366$) is considerably lower than OLO_ward ($0.573$) and SPIN ($0.542$) and ranks sixth-to-last among the $45$ evaluated algorithms. Therefore, we exclude it from our discussion.
VI-B Building a Deep Scoring Model
The accuracy of our scoring method in measuring the quality of visual patterns makes it an appealing optimization criterion for reordering algorithms. However, this method requires prior knowledge of the types and sizes of the patterns to configure the convolutional kernels. This prerequisite, often absent in matrices beyond this benchmark, limits the generalizability of the scoring method. To address this, we build a unified scoring model based on ReorderBench. This model aligns with the convolution- and entropy-based scoring method across all four visual patterns in both binary and continuous matrices and can also measure matrices of varying sizes. As shown in Fig. 14, the key feature of this model is that it does not require prior knowledge of the patterns in the matrix, which greatly improves its generalizability to matrices beyond ReorderBench. Next, we introduce how to build this model.
@@ VI-B Building a Deep Scoring Model
Model training. Each ReorderBench sample includes: 1. matrix $A_{i}$, 2. the type of pattern in its template $pt_{i}$, 3. the ground-truth quality score $S_{i}$, and 4. the number of index swaps applied to the matrix $sw_{i}$. A key characteristic of matrices with high-quality patterns is that similar rows are adjacent. Existing scoring methods check this adjacency by taking the dissimilarity matrix as input [39]. Consequently, we augment each sample with dissimilarity matrix $D_{i}$. In line with common practice in matrix reordering [2, 20], we derive the dissimilarity matrix $D_{i}$ from $A_{i}$ using the Euclidean distance between its rows. This method is chosen for its simplicity, computational efficiency, and proven effectiveness in capturing the differences between rows in a matrix.
@@
Scoring matrices of varying sizes. To predict scores for matrices of varying sizes, the key is transforming the matrix to the desired size while preserving visual patterns. Previous research in computer vision has shown that resizing an image effectively achieves this goal and incorporates it as a routine step in image pre-processing [76]. Therefore, to score matrices of different sizes, we convert them to images and resize them to the training size of the model. We resize the matrices using OpenCV [77] with the INTER__\__AREA interpolation option, as this option best preserves the visual patterns based on our detailed examination of $700$ resized matrices. The score is then predicted by the model based on the resized images. During training and evaluation, we resize all matrices to $200{\mkern-2.0mu\times\mkern-2.0mu}200$ as an example to demonstrate the effectiveness of our unified scoring model. We have verified that for $100$ larger matrices from the Network Repository with sizes up to $2000\times 2000$, downsampling to $200\times 200$ effectively preserves major visual patterns in all of them. For tasks involving even larger matrices, a new scoring model with a higher input resolution can be easily trained. For results on training scoring models with higher input resolutions, please refer to Sec. 12 of the supplemental material.
@@
We evaluate the three scoring models on matrices of all four sizes. Table IV shows their average performance. The full results are available in Sec. 11 of the supplemental material. On all matrix sizes, ConvNeXt-T consistently performs the best in terms of both accuracy and efficiency. Therefore, we employ ConvNeXt-T to build the unified scoring model. We demonstrate the capability of the scoring model to handle real-world matrices by applying it to the Pajek graph collection [8] and the Network Repository [12]. To align with the training samples, we normalize these matrices so that their entries fall within $[0,1]$. As shown in Fig. 15, the unified scoring model generates reasonable scores. For example, the first matrix contains a high-quality block pattern with a score of $0.73$. This score represents the percentage of the identified block pattern relative to the inherent ones in this matrix.
@@
VI-C Building a Matrix-Reordering Model
The representative and diverse matrices in ReorderBench offer valuable supervision for training deep reordering models. Existing deep learning-based reordering methods are typically trained on limited datasets that consist of either the given matrix [3] or its reordering results [4]. These training methods aim to learn features specific to the given matrix, resulting in models that are adept at reordering only the given matrix and thus lack generalizability. To address this limitation, we treat the index-swapped matrices as negative samples and their corresponding ground-truth matrices as positive samples. This strategy allows our model to learn a broader range of features, facilitating the ability to generalize across previously unseen matrices. The model is based on ResNet [71] due to its demonstrated performance. The model architecture is introduced in Sec. 13 of the supplemental material.
@@ VI-C Building a Matrix-Reordering Model
We train eight deep reordering models for each matrix size, with four dedicated to binary matrices and four to continuous matrices. Within each matrix type, four models correspond to four different visual patterns. All these models are trained with a batch size of 512 using AdamW optimizer with a base learning rate of 0.01. For each model, we train for 120 epochs and employ a cosine annealing scheduler with 20 epochs of linear warm-up. The models trained on each matrix type are integrated into an ensemble-based method [78], where the unified scoring model selects the best reordering result from these models. We provide more details on the ensemble-based method in Sec. 9.4 of the supplemental material. To further enhance the model performance and robustness, we perform test-time augmentation based on the unified scoring model. More details on the augmentation are described in Sec. 14 of the supplemental material.
Results. The last row of Table III presents the performance of our deep reordering model averaged over all matrix sizes. The full results are available in Sec. 11 of the supplemental material. For the block pattern, our model performs slightly better than existing methods. As a central topic for matrix-reordering research, the task of revealing block patterns is well studied and poses a challenge for further improvement. For the off-diagonal block, star, and band patterns, our deep reordering model achieves higher performance scores than the existing methods. Although ResNet-18 may seem basic by current standards, the observed improvement confirms the value of our benchmark. Moreover, although reordering continuous matrices is generally harder than binary matrices, our model achieves higher performance scores for all four visual patterns in continuous matrices due to the larger amount of training data. This also highlights the importance of large-scale datasets for training high-performance deep reordering models.
@@ VII Discussion and Future Work
Support hybrid patterns. Although the proposed pipeline achieves impressive results in generating matrices with the four visual patterns, the creation of matrices with hybrid patterns, such as the combination of block + band, poses two challenges. The first is how to generate representative matrices with hybrid patterns. Addressing this involves exploring methods to integrate different patterns into hybrid ones, which deserves further investigation. A promising method is to adapt the mix-up technique [79], commonly used to augment training data. This technique improves the robustness and generalization of the model by blending multiple images and their labels to create new composite samples. The second lies in assessing hybrid patterns. As different types of visual patterns in a matrix can potentially obscure each other, and new patterns may appear during variation generation, it is difficult to measure the quality of these hybrid patterns by simply combining their scores derived by our scoring method. For example, Fig.  16(a) shows a matrix with a hybrid pattern of off-diagonal block and band. After 30 times of index swaps, these two visual patterns obscure each other, and many star patterns also appear (Fig.  16(b)). This highlights the importance of exploring effective methods to identify and measure the quality of hybrid patterns when they are mixed during the variation process. Although our initial attempts to handle hybrid patterns, presented in Sec. 9 of the supplemental material, are promising, fully addressing these challenges will require significant further effort and deserves a separate publication.
@@
Reorder matrices of varying sizes with one model. Our evaluation demonstrates that, compared with existing algorithms, deep models greatly improve reordering performance by predicting the best permutation instead of performing a step-by-step search. However, the adaptability of deep reordering models to matrices with varying sizes remains an issue. These models are generally designed for and trained on matrices of fixed size [4, 3]. This highlights a potential area for innovation, developing foundation models to accommodate matrices of different sizes [80]. For smaller matrices, padding the matrix with empty rows and columns allows it to be reordered by the deep model. However, trials on our reordering model show that it occasionally ignores the matrix with smaller patterns, as the addition of many empty rows and columns makes these patterns nearly indistinguishable from noise. Thus, developing techniques to preserve original patterns during padding warrants further exploration. For larger matrices, one possible solution is to employ the divide-and-conquer strategy. The matrix is first divided into smaller matrices, each of which is reordered independently. Then they are combined into the final reordered matrix. However, how to divide the matrix without damaging the visual patterns remains an issue, which deserves further exploration.
@@ VIII Conclusion
In this paper, we introduce ReorderBench, a benchmark designed to evaluate and improve matrix-reordering techniques. It features a representative and diverse collection of binary and continuous matrices, and incorporates a convolution- and entropy-based scoring method. Three applications have demonstrated that this benchmark not only serves as a resource for evaluating different reordering algorithms but can also enable more effective and efficient development of a unified scoring model and deep reordering models. These improvements enhance our ability to reveal and interpret the underlying structures within complex datasets.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under grant U21A20469 and in part by the Tsinghua University-China Telecom Wanwei Joint Research Center. The authors would like to thank Duan Li, Zhen Li, and Yukai Guo for their valuable discussions and comments.